### v1

# The Impact of Large Language Models (LLMs) on Code Review Process

Abstract
[Context:] Large language models (LLMs) have recently gained prominence in the field of software development, significantly boosting productivity and simplifying teamwork. Although prior studies have examined task-specific applications, the phase-specific effects of LLM assistance on the efficiency of code review processes remain underexplored. [Objectives:] This research investigates the effect of GPT on GitHub pull request (PR) workflows, with a focus on reducing resolution time, optimizing phase-specific performance, and assisting developers. [Methods:] We curated a dataset of 25,473 PRs from 9,254 GitHub projects and identified GPT-assisted PRs using a semi-automated heuristic approach that combines keyword-based detection, regular expression filtering, and manual verification until achieving 95% labeling accuracy. We then applied statistical modeling, including multiple linear regression and Mann–Whitney U test, to evaluate differences between GPT-assisted and non-assisted PRs, both at the overall resolution level and across distinct review phases. [Results:] Our research has revealed that early adoption of GPT can substantially boost the effectiveness of the PR process, leading to considerable time savings at various stages. Our findings suggest that GPT-assisted PRs reduced median resolution time by more than 60% (9 hours compared to 23 hours for non-assisted PRs). We discovered that utilizing GPT can reduce the review time by 33%, and the waiting time before acceptance by 87%. Analyzing a sample dataset of 300 GPT-assisted PRs, we discovered that developers predominantly use GPT for code optimization (60%), bug fixing (26%), and documentation updates (12%). [Conclusion:] Our findings highlight the potential of GPT to accelerate code reviews, improve code quality, and address persistent bottlenecks. This research sheds light on the impact of the GPT model on the code review process, offering actionable insights for software teams seeking to enhance workflows and promote seamless collaboration.
keywords:
LLMs , GPT , AI assistance , pull requests , pull request phases , code review process , developer productivity.
journal: Information and Software Technology
[…]
1 Introduction
Modern software development is heavily based on collaborative workflows, with Pull Requests (PR) serving as a central mechanism to integrate code changes [1, 2, 3]. A PR goes through multiple phases, including submission, review, change, and merging, each requiring coordination between contributors and reviewers [1]. Effective PR management is crucial to maintaining software quality, but it remains a bottleneck in many projects due to long review times, inconsistent feedback, and inefficient communication [4]. Developers often experience delays due to unclear review comments, excessive revisions, and prolonged waiting times before changes are accepted [5].
Inefficiencies in the software development cycle, such as long review times, can lead to frustration among team members, decreased productivity, and even project delays [6]. As codebases become increasingly intricate, there is a growing need for sophisticated automation tools to facilitate the code review process and improve collaboration and teamwork [7]. In parallel, LLMs, such as Generative Pre-trained Transformers (GPTs) [8] have rapidly evolved, demonstrating strong capabilities in code generation, bug detection, and documentation assistance [9, 10]. These models are capable of suggesting enhancements to existing code, resolving unclear logical structures, and automating repetitive tasks, thereby becoming indispensable tools for software engineers.
Despite the increasing adoption of LLM-powered tools in development workflows, a systematic investigation into their impact on PR processes is lacking. Previous research on the use of LLM in the PR context [9, 11, 12, 13] has focused mainly on task-specific applications, such as commit message generation [14] and documentation support [15], while neglecting to explore their phase-specific impact across the entire PR lifecycle. Furthermore, while LLM-generated suggestions can be beneficial, it is necessary to quantify their effectiveness in improving review speed and reducing PR merge time.
This study aims to fill this knowledge gap by empirically evaluating the impact of GPT on code review waiting times, with a particular focus on phase-specific effects. It examines their influence on resolution time, their role in different PR phases, and the reasons why contributors and reviewers adopt GPT as assistance. Given the proven effectiveness of GPT on a wide range of software engineering tasks, including code generation, debugging, and documentation, understanding how and to what extent GPT enhances the effectiveness of PRs in each phase is crucial to assess their usefulness in real-world operations. These insights can guide the design of more efficient GPT-supported code review workflows and inform best practices for integrating AI assistance into collaborative software development.
Toward this goal, we examine the PRs that software developers actively engaged with GPT to seek recommendations for code review tasks or problem-solving. We refer to these as GPT-assisted PRs throughout the paper. In contrast, GPT non-assisted PRs are those where developers either do not use GPT at all or use it strictly for implementation tasks (i.e., making an API call to a GPT model), without relying on it for code review insights or decision making. Our research investigates how GPT-assisted PRs influence the duration and quality of the code review process. For this purpose, we start by curating a dataset of 25,473 PRs from 9,254 projects on GitHub. Next, we extract six features (project_age, no_commits, files_added, files_deleted, files_changed, time_to_merge) to characterize projects, PRs, and code reviews. We then analyze how these factors influence the review and resolution times [16]. Finally, we also investigate which development tasks involve the use of GPT-assisted PRs throughout the PR phases, aiming to identify where such assistance is most effectively integrated. In summary, our aim is to answer the following three research questions:
  1. RQ1:
    Does GPT-assistance during the PR process reduce the resolution time? We aim to understand the potential of LLMs to expedite the resolution of PRs in the code review process. Specifically, we analyze their impact on the total time required to resolve a PR, from its submission to completion. Our results suggest a correlation between the use of GPT-assisted PRs and a decrease in merge time approximately from 23 to 9 hours.
  2. RQ2:
    How does GPT assistance impact various phases of the PR process? We aim to further explore the effect of GPT in various phases of the PR process, such as submission, evaluation, modification, and waiting time, to determine where its support is most advantageous. Our analysis shows that GPT-assisted PRs complete faster across key phases of the review process. Specifically, they take 1 hour at the Review phase, 3 hours at Waiting for Changes, and 1 hour at Change, compared to 3 hours, 24 hours, and 1 hour, respectively, for non-assisted PRs. These results highlight the potential of GPT to streamline the PR workflow, especially by reducing waiting times.
  3. RQ3:
    Why do contributors and reviewers use GPT during the PR process? We aim to better understand how programmers use GPT to enhance code quality, fix errors, and improve documentation. By analyzing these application patterns, we uncover how developers integrate this technology into their daily workflows, highlighting its most potential benefits. Our findings show that developers devote a significant part of their review time to refining and correcting the source code. Specifically, 60.26% of enhancement tasks occur during the Review phase, 42.31% during Waiting for Changes, and 35.90% during the Change phase. This explains the role of GPT as assistance in enhancing efficiency for those time-consuming activities.
Finally, our research provides recommendations for software developers, both maintainers and contributors, as well as researchers, aimed at optimizing and accelerating the PR review process. We argue that applying the GPT recommendations as a support tool can strengthen collaboration by allowing participants to better anticipate and resolve potential issues.
Our Contributions. In summary, we make the following contributions in this paper:
  • To the best of our knowledge, this is the first comprehensive analysis that examines the effectiveness of GPT during the PR review phases.
  • We present empirical evidence on the impact of GPT on PR-based development phases and identify the types of software development activities where it is used the most frequently.
  • To promote the reproducibility of our study and facilitate future research on this topic, we publicly share our scripts and dataset online at https://github.com/acollant/GPT-Assistance-PR.
Paper Organization. The remainder of the paper is organized as follows. We review related work in Section 2 and describe our study setup and dataset in Section 3. We report our approach and findings for each research question in Section 4 and discuss the implications of our findings in Section 5. Finally, we describe the limitations of our study in Section 6 and conclude the paper in Section 7.
2 Related Work
In the following, we first review research on the challenges in code reviews and collaboration before moving on to an overview of factors affecting PR lifetime and latency. Finally, we review studies on the application of LLMs in code reviews.
Challenges in Modern Code Reviews (MCR) and Collaboration. Code reviews are a critical aspect of modern software development, ensuring code quality, maintainability, and adherence to best practices. However, the process comes with several challenges that can affect efficiency, collaboration, and overall development workflow. For example, long waiting times where PRs often remain unreviewed for extended periods, slowing down development. Kudrjavets et al. [5] reported that between 29% and 63% of the overall code review lifetime is spent waiting for the accepted changes to merge. Bacchelli and Bird [3] reveal that scheduling and time issues also appeared challenging; yet, understanding the code and comments takes most of the review time. According to these authors, code reviews tend to be less focused on detecting defects than expected and, instead, provide additional advantages, such as knowledge transfer, increased team awareness, and creation of alternative solutions to problems. Davila and Nunes [17] also reported that reviewers have found it difficult to understand the code, which prompts the need for enhancements in visualizing code changes to facilitate understanding. However, a recent study has shown that ChatGPT is an alternative to generating some kind of explanations during MCR [18]. Davila et al. [15] identified five challenges to adopting generative AI in MCR: trustworthiness, lack of context, misleading review comments, security and privacy, and token limits. Although prior studies have explored aspects of the use of GPT in software development, they do not provide a systematic analysis of its impact on waiting times and efficiency in all phases of the PR workflow. Our study addresses this gap by focusing on its influence during specific code review phases.
Factors Affecting PR Lifetime and Latency. Studying the factors that influence PR latency is crucial for improving software development efficiency, collaboration, and overall code quality. Understanding what causes delays in PR review and merging can lead to optimizations that benefit both developers and organizations [4], [19], [16], [20], and [21]. Zhang et al. [4] identified that the number of commits at open time is negatively correlated with PR merging but shifts to a positive correlation at close time. Rose [19] built a Multiple Linear Regression Model (MLR) that indicates that the PR size has a statistically significant effect on the review time, as the positive value of the regression coefficient implies that as the PR size increases, the review time for PRs also increases. Moreira Soares et al. [16] further emphasized that PRs possess some structural characteristics, such as the number of modified files, the number of commits, and the number of changed lines of code. Moreira Soares et al. [16] concluded that the combination of structural characteristics may increase the chances of PRs having a very short and long lifetime.
Although prior research explores factors affecting PR latency (e.g., size, number of commits), there is limited research into how GPT can have an effect on PR resolution time. Khatoonabadi et al. [20] developed a machine learning model to predict the latency of the initial responses of the maintainers and contributors after receiving the first feedback from the maintainer. These authors highlighted that the timing of submissions and feedback are factors that have a significant influence on response waiting times. Their approach, which provides estimated waiting times, helps open-source projects collaborate better by allowing maintainers and contributors to anticipate and address potential delays in PR reviews. This study complements the aforementioned work by offering empirical insights into GPT usage beyond commit messages, specifically examining detailed time reductions in key phases such as review and change.
The Application of LLMs in Code Reviews. LLM-powered tools can offer services designed to streamline various code review-related development tasks. Research has been conducted on the impact of AI-driven generative tools in this context [9], [14], [15], [13], [22], [23], [24], [25]. For instance, Davila et al. [15] conducted a literature review on the application of AI-based generative tools in a code review context and highlighted that AI-based tools can positively impact code review efficiency by reducing reviewer workload and review duration and increasing the likelihood of acceptance in GitHub projects. Xiao et al. [14] observed that PRs assisted by Copilot have a 1.57 times higher likelihood of being merged compared to those not assisted by Copilot for PR. Tufano et al. [9] reported how developers use ChatGPT to automatically generate a commit message and a PR/issue description. Rasheed et al. [13] highlighted the potential of LLM-powered AI agents to considerably improve conventional code review procedures. These agents provide a twofold advantage: they enhance code quality and improve developer learning.
Vijayvergiya et al. [22] shared the experience in developing, implementing, and evaluation AutoCommenter, a LLM-powered code review tool that enforces coding best practices and has a positive impact on the developer’s workflow. Yu et al. [23] proposed Carllm (Comprehensibility of Automated Code Review using LLMs) that improves the comprehensibility of automated code review. Hao et al. [24] studied the role of ChatGPT in collaborative coding by analyzing the shared conversation between developers and ChatGPT in GitHub PRs and issues, finding that the most frequent requests involve code generation, conceptual understanding, how-to guidance, issue resolution, and code review. Studies on GPT have examined particular tasks (e.g., code generation, commit message generation, or bug fixing) but lack a holistic view of its phase-specific impacts during the PR life cycle. Our research fills this void by examining the use and impact of GPT on specific phases, including the review process and the waiting period.
3 Study Setup
Figure 1: An overview of our study setup that describes a semi-automated approach for identifying, detecting, and mapping GPT-assisted PRs to the phases in which GPT is used
In this paper, we investigate the impact of using GPT as an assistant tool in the PR code review process. Specifically, we examine how GPT influences the overall PR resolution time and how its effects vary between different phases of the review workflow. We also aim to identify the types of review tasks for which developers actively seek GPT support. Figure 1 presents an overview of our study design. We begin with an automated procedure to identify potential GPT-assisted PRs, followed by a semi-automated verification step to confirm the actual use of GPT during the review. Finally, we map the verified GPT-assisted PRs to their corresponding review phases. The remainder of this section provides a detailed explanation of each step in our methodology.
3.1 Searching Potential GPT-assisted PRs
To identify relevant PRs for our analysis, we rely on GitHub, which adopts a PR-based approach to submit, review, and merge code contributions [17]. We use the GitHub REST API [26] to retrieve a broad set of candidate PRs in which the assistant of a GPT model could have been used during the review process.
3.1.1 Searching
Following a similar strategy to that of  Tufano et al. [9], we conduct keyword-based searches for mentions of ”GPT” in four main PR fields.
  1. $\blacksquare$
    Project Name: Although some GPT-focused projects may not actually use GPT for PR reviews, we still record these mentions before filtering out irrelevant cases [27].
  2. $\blacksquare$
    PR Title: The use of GPT may be evident in the PR title (e.g., refactored using GPT suggestions [28]).
  3. $\blacksquare$
    Body+Comments: A PR description and/or comments can refer to GPT explicitly (e.g., code generated by GPT [29]).
  4. $\blacksquare$
    File(s) Name: References in filenames (e.g., chatgpt_review.md [30]) can indicate GPT-generated artifacts.
3.1.2 Filtering
To avoid projects with minimal activity or that lack meaningful collaboration, we keep only projects that meet the following criteria: (1) the repository of the PR has at least 10 stars [31], to remove toys or trivial projects. (2) at least two contributors are involved in the review process, to exclude self-reviewed PRs.
3.1.3 Downloading
Based on the aforementioned criteria, a total of 25,473 pull requests were collected from 9,254 repositories on May 9, 2024. For each PR, we extracted detailed metadata, including the associated project name, URL, PR number and title, event history, contributor information, creation date, comments, and the set of modified files.
3.2 Identifying GPT-assisted PRs
In this semi-automated step, our goal is to define the heuristics that indicate the use of GPT to classify PRs as GPT-assisted or GPT-non-assisted. By doing this, we exclude false positives (i.e., PRs that mentioned GPT as discussed in Section 3.1 but did not use the GPT model to assist in the code review process). This step ensures that the PRs used for our analysis are PRs where GPT models are used in the code review process, rather than just being part of the project code.
3.2.1 Random Samples
With the goal of discovering patterns related to the use of GPT in PRs, we randomly sampled a total of 450 GitHub PRs to examine indicators of GPT assistance in code review activities. Three authors independently inspected 150 unique PRs in three separate rounds, analyzing project names, PR titles, modified files, and PR bodies. In Round 1, we reviewed the first set of 150 PRs and cross-validated our observations to ensure that the identified patterns consistently reflected GPT-assisted contributions. In Rounds 2 and 3, we applied the same procedure to two additional sets of 150 PRs each. After each round, the reviewers met to consolidate recurring patterns, clarify ambiguous cases, and resolve classification disagreements through discussion until consensus was reached. The findings of Round 1 are presented in Table 1. In particular, most PRs were ultimately classified as GPT-non-assisted, as GPT was often integrated into the software (e.g., through GPT APIs) but not explicitly used to support the code review itself.
3.2.2 Defining Heuristics
We defined four heuristics to analyze and label the presence of GPT-assisted PRs, as outlined in Table 2. These heuristics are reinforced by a regular expression-based technique (detailed in the appendix), which automatically classifies the collected PR instances as either GPT-assisted or GPT-non-assisted based on the identified patterns: H1: Project name, H2: PR title, H3: Modified files, H4: PR body+comments.
3.2.3 Applying Heuristics
In the initial phase of our analysis, we experimented with various sizes of n-gram to determine which configurations offered the most valuable insights for detecting recurring patterns. In contrast to the methodology employed by  Tufano et al. [9], which used 2- or 3-grams, we found that a maximum size of 5-n-gram provided sufficient context or relevant information to accurately capture the role of GPT in GPT-assisted pull requests. We then proceeded to automatically extract both forward and backward n-grams, up to a maximum of 5-grams [32], containing the term GPT, while simultaneously applying a set of heuristics designed to reduce the occurrence of false positives. For instance, when analyzing the PR body, we encountered the sentence: ”Instead of incorporating muP into GPT-NeoX we are going to move these changes to [our fork of their repo]” [^1] https://github.com/EleutherAI/gpt-neox/pull/1061#issuecomment-1771635901. From this, we extracted the backward n-gram ”of incorporating muP into GPT-NeoX” and the forward n-gram ”GPT-NeoX we are going to”; the application of heuristic H4 identified GPT-NeoX as a reference to the PR name, thereby classifying this PR as non-GPT-assisted. Additionally, we applied heuristics to filter and automatically eliminate false positives, including cases where GPT is mentioned in a PR but not used for assistance, as well as instances where a PR may appear GPT-assisted but is actually unrelated (e.g., mentions of GPT in the context of non-code-related discussions, such as research or tool integration). These measures ensured that only relevant PRs, where GPT is actively employed for code review assistance, were included in the analysis.
3.2.4 Inspection
To ensure the reliability of our heuristics, two authors independently examined 200 randomly chosen PRs by carefully reviewing the labels generated during the execution of the heuristics. They then discussed unambiguous cases where the labeling was unclear to reach an agreement, such as whether the PR used GPT as an assistant. In the final step, the heuristics were adjusted and reapplied to improve accuracy and consistency. We performed three rounds of refinement/reapply until we obtained an accuracy of 95% as an acceptable agreement score in the labeling of PRs.
Table 1: GPT PRs During a Round of Sampling Analysis
Filter Assistance Non Assistance
Project name 4 21
PR title 6 19
Modified files 23 58
PR body 1 24
Total 34 122
Note: Some GPT-assisted PRs matched two or more filters simultaneously. The total count (34) reflects overlaps between categories, not distinct PRs per filter. For example, a single PR might be identified as GPT-assisted based on both its title and modified files, contributing to multiple row counts.
3.2.5 Combining the Heuristics Outputs
We reapplied the heuristics to the entire dataset to discard a higher number of false positives. All PRs were labelled GPT-assisted or GPT-non-assisted after applying the heuristics (as detailed in Table 2). Subsequently, one of the authors examined the dataset for true positive cases by manually reviewing the final labels. Ultimately, we split the dataset into two separate groups: “GPT-assisted” and “GPT-non-assistance” PRs. The table 3 displays the distribution of GPT-assisted and GPT-non-assisted PR in the dataset.
Table 2: Definition of the Heuristics Applied to the Project Name, PR Title, PR Body, and Modified File to Identify PRs using GPT as Assistance
Heuristic Description
H1 Aims to find and filter any GPT-related keywords within the project name (e.g., EleutherAI/gpt-neo)
H2 Matches and filters any pattern that implies the use of GPT as non-assistance in the PR title (e.g., use of chatgpt-retrieval-plugin)
H3 Finds and filters any file name pattern, directory, URL, and file version that indicates the use of GPT for implementation (e.g., Add ChatGPT Image Unlocker V2, from 3.1.0 to 3.3.0) with the modified file names.
H4 Extracts and filters the patterns in H1, H2, H3 in the PR body, but including action verbs such as add, refactor, implement, and rename (e.g., Added some simple error handling around the OpenAI ChatGPT logic to capture rate limit errors and bad gateways).
Table 3: Final Dataset: Assistance and Non-Assistance GPT PR Status Summary
Status Assistance Non Assistance
Open 54 38
Closed 179 219
Merged 1367 1036
Total 1600 1293
3.3 Mapping GPT-assisted to Corresponding PR Phases
In this automatic step, our goal is to define the different phases of the code review process as previously performed by Kudrjavets et al. [5]. We report descriptive statistics to quantify the extent of GPT usage in each phase, depending on whether non-productive phases are present in the review workflow.
3.3.1 Identifying PR Events using GPT
For each PR previously labeled as GPT-assisted, we applied the heuristic H4 (explained in Table 2) to timeline events [33] to label the events that the contributor and reviewers shared as GPT-assisted or GPT-non-assisted events. This specifies when and in what capacity GPT helps with PR discussions. Figure 2 is a visual representation of our definition of the PR phases:
  1. $\blacksquare$
    at_submission: the time in hours from the first commit (committed state) to when the author submits the code change (pulled state). According to figure 2, this phase corresponds to the time taken from $at\_commit_{event=0}$ to at_before_pullat\_before\_pull.
  2. $\blacksquare$
    at_review: is the amount of time, expressed in hours, that occurs between the submission of a code review (pulled state) and the first review, comment, inline comment (comment on specific code fragments), commit-commented, review-commented, provided by a person who is not the code’s author. The review phase is comparable to the time-to-first-response, which is a proportion of the time needed to accomplish the first iteration of the code review. According to figure 2, this phase comprises the period between from $at\_pull_{event!=0}$ to $at\_review_{n=1}$.
  3. $\blacksquare$
    at_change: refers to the time in hour from the last review, comment, inline comment (comment on specific code fragments), commit-commented, review-commented until the acceptance of the code review by someone other than the author of the code. According to figure 2, this phase covers the time between $at\_review_{n}$ and at_acceptedat\_accepted.
  4. $\blacksquare$
    at_waiting_time: the time in hours spent on multi-turn conversations and feedback shared between reviewers and the author. The first waiting time is between $at\_review_{2}$ and $at\_reviewn_{n-1}$. GPT was not use for the second waiting time phase between acceptance and merge/close.
Figure 2: Pull Request Phases
To analyze the impact of LLMs on each PR phase, we mapped a chronological sequence of events (including pulling, examining, integrating, forcing the head reference, commenting on reviews, comments in lines, commit messages, and merging) to the phases. To guarantee the accuracy of our findings, we only considered PRs with sequential timestamps. In these cases, the review phase follows the submission phase, the approval step follows the review, and the merging occurs after acceptance. Finally, to ensure human participation in the review process, we excluded PRs in which the only participants, either as contributors or reviewers, were the original author and automated bots.
To have a fair comparison between the GPT-assisted and GPT-non-assisted PRs, we quantify the similarity between them by applying the Manhattan [34, 35] on the same factors as applied in RQ1 4.1. To mitigate the presence of outliers or bias in the factor, we rescale the no_commits, PR_size, no_changed_files, project_age to a range within [0, 1], applying the max-min method [35]. Since the two well-known metrics (the Manhattan and Euclidean distance produce very similar results (Spearman correlation $\rho=0.98$ [36] when computing the distances between factors), we finally test the differences between the PRs supported by or without GPT by getting the minimum distance value from the Manhattan distance.
4 Results
4.1 Does GPT-assistance during the PR process reduce the resolution time?
Motivation: Recent studies have investigated the influence and popularity of ChatGPT on code review [37], [11], [24], [38]. The potential impact of using GPT models as assistants throughout the entire PR lifecycle remains largely unexplored. This research question investigates whether relying on GPT-assisted PRs affects code review resolution time in GitHub projects by analyzing the impact of GPT models on the overall merge time of PRs.
Approach: To answer this research question, we build a Multiple Linear Regression Model (MLRM) [39] to empirically explain PR lifetime. For this purpose, we examine previous studies to identify a list of known factors that significantly influence PR resolution time [4], [16], [19], [40]. Bernardo et al. [41] defines resolution time as the time it takes for a code change to be merged into the development branch after review. This implies that the change is not fully implemented until it can be built, tested, and deployed. This is possible only after merging into the main branch. We will use the time it takes for code modifications to be merged as a metric, selecting only PRs that have been merged, considering the identified factors that affect this time [40]. Furthermore, we added a two-state indicator to the dataset, with 1 representing PR that use GPT-assisted and 0 representing PR that use GPT non-assisted. This indicator is used to better understand the impact of the GPT as a factor on PR latency.
After this step, our dataset includes only the merged PRs with a total of 1,367 PRs (85.43% of the total GPT-assisted PR population as shown in Table 3) using GPT-assisted. To compare the resolution time between the GPT-assisted PRs and GPT-non-assisted PRs, we searched for the most similar PRs among the 23,874 identified GPT-non-assisted PRs by using the Manhattan distance [42] as shown in Figure 1. To compare the statistical difference between PRs using GPT-assisted or GPT-non-assisted, as an influencing factor, we use the Mann-Whitney U test, a non-parametric test that compares two independent groups that do not require normally distributed samples[43]. To carry out this statistical test, we utilize the stats package [^2] R Core Team. 2023. R: a language and environment for statistical computing. https://www.R-project.org considering a confidence level of 95% ($\alpha=0.05$). The following formula and list of factors characterize the Linear Model (LM) performed in our study:
$time\_to\_merge$ $=is\_gpt\_assisted$ (1)
$+no\_commits$
$+PR\_size$
$+no\_changed\_files$
$+project\_age$
  1. $\blacksquare$
    time_to_merge (dependent variable) is a numeric value that indicates the latency of PR in hours.
  2. $\blacksquare$
    is_gpt_assisted (Treatment) is a binary value ([0,1]) that translates whether a PR is a GPT-assisted PR.
  3. $\blacksquare$
    Common factors that influence the PR latency [4, 19, 16].
    1. $\square$
      no_commits (control variable) represents the number of changes pushed to a repository over time.
    2. $\square$
      PR_size (control variable) is the amount of change in the PR (added ++ deleted lines of code).
    3. $\square$
      no_changed_files (control variable) is the number of files that have changed during the review process.
  4. $\blacksquare$
    The project_age in months (control variable) variable, which represents the maturity level of the projects.
To mitigate the influence of skewness and enhance the clarity of our statistical examination, we implemented a logarithmic transformation on variables that exhibit a skewed distribution [44]. To implement our linear model shown in Equation 1, we follow the approach proposed by James et al. [39]. First, we build our model using the lm function that comes standard with the distribution [^3] https://www.rdocumentation.org/packages/stats/versions/3.6.2/topics/lm. Then we evaluate the degree of confidence of our model by obtaining the marginal and conditional coefficients of determination $(R^{2})$ [45]. To determine whether the adoption of GPT-assisted has a potential impact on PR latency, we check the statistical significance and the total variance of is_gpt-assisted in the model. To identify the level of significance of the variables, we rely on the traditional confidence level of $\alpha=0.05$ [46].
Findings: As shown in Table 4, the multiple linear regression model suggests that GPT assistance has a statistically significant effect on pull request (PR) latency. Specifically, the variable is_gpt-assisted yields a p-value of $4.88e^{-11}$, well below the standard significance threshold of $\alpha=0.05$, indicating strong evidence against the null hypothesis. The corresponding negative coefficient suggests that GPT assistance is associated with a reduction in the time required to merge a PR. The standard error for this estimate is 0.0745, indicating a high degree of precision, and the large associated t-value reinforces the robustness of the finding. Together, these results demonstrate that GPT-assisted PRs tend to be merged more quickly than their non-assisted counterparts. This quantitative result is consistent with the descriptive statistics presented in Figure 3, where GPT-assisted PRs have a median merge time of approximately 9 hours, compared to 23 hours for non-assisted PRs.
Table 4: Linear Model Coefficients for log1p(time_to_merge_h)
Predictor Estimate Std.Err t value Pr(>|>| t $|$) Signif. (Intercept) 0.5186 0.1248 4.154 3.37e-05 *** is_gpt-assisted -0.4921 0.0745 -6.601 4.88e-11 *** log1p(no_commits) 0.8052 0.0715 11.267 << 2e-16 *** log1p(PR_size) 0.1576 0.0322 4.899 1.02e-06 *** log1p(no_changed_files) 0.0064 0.0702 0.091 0.927 log1p(project_age_month) 0.3371 0.0275 12.252 << 2e-16 *** Signif. codes: 0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1
Figure 3: Distribution of time to merge for GPT-assisted and GPT non-assisted PRs
Answer to RQ1. The results highlight the potential of GPT assistance to significantly accelerate the pull request review process, reducing merge times by more than 60%. Integrating GPT as an assistance in the development workflows can help streamline code review, improve team efficiency, and facilitate faster software delivery.
4.2 How does GPT assistance impact various phases of the PR process?
Motivation: Previous studies have shown that factors such as technology, collaboration dynamics, and project context influence the effectiveness of code reviews [4, 5, 6, 41]. More recently, AI-driven tools, particularly LLMs, have been explored for their potential to assist with PR-related development tasks and improve review efficiency [15, 14, 9]. Despite growing interest, little is known about how developers actually use GPT during the code review process or how its usage varies across different phases of a PR lifecycle. This gap motivates our investigation of the impact of GPT as assistance on review and resolution time. By identifying the specific phases in which GPT is applied and understanding its role as an assistive (rather than an automated) tool, we aim to assess where it offers the greatest benefit in the review workflow.
Approach: To investigate the impact of using GPT as assistance throughout the PR phases, we first define the distinct phases of the PR workflow, starting from the submission of a change to merge into the project branch including the presence of the waiting time, as described in detail in the section  3.3. To assess whether the use of GPT-assisted leads to statistically significant differences in PR phases, the Mann-Whitney U test was applied [47]. This non-parametric test is well-suited for comparing two independent samples, particularly when the assumption of normality cannot be guaranteed. In this study, the GPT-assisted and GPT-non-assisted datasets are treated as independent groups, as each PR belongs to only one condition (each PR is placed into only one of these two groups), with no overlapping or paired observations. Using the Mann-Whitney U test aims to determine whether the distributions of phases differ significantly between the two groups, thus providing evidence on the potential impact of GPT-assisted.
Findings: The results presented in Table 5 highlight significant differences in the duration of review between GPT-assisted and non-assisted PRs, with particularly notable reductions observed during the At Review and At Waiting Before Change phases. In the At Review phase, the median time for GPT-assisted PRs was approximately 1 hour, compared to 3 hours for GPT-non-assisted PRs — reflecting a 66.7% reduction in the review time. The most significant difference was found in the At Waiting Before Change phase, where GPT-assisted PRs had a median waiting time of approximately 3 hours, while their non-assisted counterparts experienced a median of 24 hours. This corresponds to an 87.5% reduction in waiting time. The results of the Mann–Whitney U test confirmed that the reductions in these two phases were statistically significant, with p-values well below the conventional significance threshold ($\alpha$ = 0.05). These low p-values provide strong evidence against the null hypothesis, suggesting that the differences are unlikely to be due to random variation. Interestingly, no significant differences were observed in Phase 1 (At Submission) and Phase 5 (At Waiting After Acceptance), where there was no clear indication of GPT as assistance was used during these stages, suggesting that the use of GPT is concentrated in the core iterative phases of the review process.
Table 5: Median time (in hours) spent across three key PR phases for GPT-assisted and non-assisted pull requests
GPT At Review At Waiting for change At Change Assisted 1 3 1 Non-Assisted 3 24 1
Note: There is no evidence of GPT assistance being used during the submission or waiting after acceptance phases of the PR workflow.
Answer to RQ2. Our analysis reveals that GPT-assisted PRs are associated with significantly shorter durations in specific phases. The review phase is 33% faster, and the revision before acceptance is reduced by 87% compared to GPT-non-assisted PRs. However, no significant differences are observed in the change phase.
4.3 Why do contributors and reviewers use GPT during the PR process?
Motivation: Our prior analysis revealed that the use of GPT was associated with notable variations in the duration of individual PR phases, suggesting that its influence extends beyond the overall resolution time to phase-specific developer activities. While earlier studies have examined how developers use GPT for specific code review tasks [9, 12], recent work has raised concerns about its influence on review discussions and developer responses [24, 38]. Although GPT tools show promise across various software engineering activities, little is known about their role across different phases of the PR workflow. This suggests an underexplored but potentially broad impact of GPT-based tools on the code review process. To fill the gap, we analyze the purposes for which developers use GPT within GitHub PR phases.
Approach: To explore the purposes for which developers use GPT as assistance, we performed a detailed analysis of a statistical sample of 310 randomly selected GPT-assisted PRs drawn from the dataset described in Section 3.3. Initially, two authors independently examined and assigned labels to each instance (e.g., enhancement, documentation, fixing bugs). The labeling process was conducted in two rounds, each covering about 155 GPT-assisted PRs. After each round, the authors met to refine the labels, and the final classification was based on the taxonomy proposed by Tufano et al. [9]. The analysis specifically targeted events in which developers explicitly recorded their prompts to GPT and the corresponding suggestions generated by the model. This classification enabled a structured understanding of the functional contexts in which GPT-assisted PRs are integrated into the PR workflow.
Findings: Table 6 presents the most common tasks for which developers seek the assistance of GPT during each phase of the PR process, along with details on their frequency and significance, highlighting its focused role in supporting diverse development tasks. In the following, we provide a detailed description of these tasks and their distribution across the different PR phases, highlighting the overall use of GPT on the code review lifecycle.
  1. $\blacksquare$
    Enhancement: As shown in Table 6, this is the most common activity during the stages (60.26% At Review, 42.31% Waiting before Change and 35.90% At Change), when developers offer current code as a baseline for adaptation. Most of the suggestions made by GPT are about refactoring the code, missing error handling, and making changes to existing code for performance reasons. We can see that the rewriting of the code was for minor modifications; for example, the most common modification was renaming a variable, function or class name, changing a flag variable, declaring a variable, adding/removing a conditional or reorganization a function. In contrast, we see more occasions when GPT proposes adding an error-handling instruction (try..catch). An example of a case of enhancement can be found in the PR [48], where the developer asked ChatGPT to change a Ruby plugin: updates external-posts.rb plugin, allowing the user to specify an explicit lists of urls in config.yml that are then displayed in the blog feed as external posts. 99% of the code in this change is written by gpt-4: https://chat.openai.com/share/24432d24-36a7-4d6f-a5c0-d7e5142f68cd
  2. $\blacksquare$
    Implementation: Contrary to our expectations, in our sample we observe that the Implementation in which the developers ask GPT to generate a piece of code (from scratch) has a low frequency, as shown in Table 6 where the ratio is split on 6.41% At Review, 10.26% Waiting before Change and 8.97% At Change. An illustrative instance is presented in [49], where a developer sought assistance from GPT in generating code to retrieve remote resources, and GPT responded with a Python-based implementation. We identified several cases in which reviewers expressed disagreement with the solutions generated by GPT. One such example is documented in [50], where the reviewer remarked, ”I will not blindly accept any AI-generated code” and ”using GPT so might not work as intended.”.
  3. $\blacksquare$
    Bug fixing:. This task ranks as the second most common use case for GPT assistance, occurring primarily during the Review phase (25.64%) and to a lesser extent during the Change phase (21.80%). In many instances, developers consult GPT to diagnose why a specific piece of code is not functioning correctly or to request help in fixing an error (e.g., “GPT found 2 config files not exist in the dataset list.”) [51]. GPT-generated suggestions also include improvements such as adding error-handling logic —for example: “The code does not seem to have any error handling. This could lead to unexpected behavior if an error occurs. Consider adding try-catch blocks to handle potential errors.” [52].
  4. $\blacksquare$
    Testing: We find few cases where developers ask for GPT-assisted about testing related tasks in our sample (every single character of tests is generated by GPT-4) [53], (I have added the shorthand option -f for the –files flag in the _ask.py file and updated the test case in test_gpt_cli.py.)[27]. We also notice that the testing inquiries to GPT were for more complex tasks such as classes and core logic testing (Chat GPT says this might work: the class AZDemoContentTest extends QuickstartFunctionalTestBase…)[54].
  5. $\blacksquare$
    Documentation: We observe that developers utilize GPT both for generating new documentation — such as implementation descriptions in markdown files or tutorials on how to use specific functionalities [55] — and for modifying existing documentation, including improving navigability, enhancing readability and formatting, and updating outdated links [56]. Notably, in most instances, GPT-generated suggestions pertain to the README (.md) file [57, 58].
  6. $\blacksquare$
    Other tasks: We notice that At Changed Phase and At Waiting before Change, non-programming related tasks count for the most requested activity with 28.20% and 17.95% respectively. Among the most common questions developers ask GPT for suggestions are about translation, file versions, and setting up CD/CI workflows (YAML files, file configuration, Docker, Kubernetes), Copyright, and User Interfaces [59, 60, 61].
Table 6: Frequency of Tasks That Developers Ask GPT for Assistance in PR Phases Within a Sample of 300 GPT PRs
Tasks At Review % At Waiting for change % At Change %
Enhancement 60.26 42.31 35.90
Implementation 6.41 10.26 8.97
Bug Fix 25.64 15.38 21.80
Testing 1.28 0.00 5.13
Documentation 11.54 12.82 10.26
Other 12.82 17.95 28.20
Note: Some GPT-assisted PRs matched two or more filters simultaneously. The total count (34) reflects overlaps between categories, not distinct PRs per filter. For example, a single PR might be identified as GPT-assisted based on both its title and modified files, contributing to multiple row counts.
Answer to RQ3. Contributors and reviewers mainly use GPT to assist with coding enhancements, bug fixes, and documentation. It is particularly useful during the review phase, helping to identify and resolve issues, refine code, and improve documentation. However, its use for creating new features or writing tests is less common. This pattern indicates that GPT primarily functions as a support tool for enhancing existing code rather than creating new components. This contributes to the effectiveness and quality of the code review process.
5 Discussion
5.1 Implications for Practitioners
These findings have important implications for practitioners. Developers can integrate GPT-assisted PR into their phases to accelerate code reviews, reduce latency, and improve collaboration. In doing so, organizations can streamline their code review processes and focus on delivering value more efficiently. In addition, establishing best practices for the use of GPT-assisted can further enhance its effectiveness and ensure seamless adoption across teams. By integrating GPT into the pull request (PR) process, teams can significantly reduce review times, expedite the implementation of changes, and minimize waiting periods, ultimately streamlining their code review cycles. This efficiency gain allows developers to focus more on delivering value rather than being hindered by prolonged code review timelines. To maximize the effectiveness of GPT, teams should consider developing best practices for its use. For example, defining specific scenarios where GPT as assistance is most beneficial, such as refactoring code, resolving bugs, and enhancing documentation, can help ensure its targeted and efficient application. Additionally, encouraging collaboration between developers and reviewers when leveraging GPT can help maintain high-quality outcomes and prevent over-reliance on automated tools.
5.2 Implications for Researchers
An important avenue for future research lies in examining the risks of user over-reliance on GPT in code review processes. While generative AI tools like ChatGPT offer remarkable potential to streamline code review phases and enhance productivity, there is a need to understand how unproductive interactions with ChatGPT might emerge and how these could undermine efficiency or decision-making. Identifying patterns of ineffective GPT use can help researchers propose strategies to optimize its integration into code review practices in real-world scenarios. Future studies could focus on designing frameworks and interfaces that guide users toward more effective use of GPT while maintaining control of their code review workflow. For instance, developing tools that provide metrics on the quality (i.e., reliability, usability, maintainability, testability of the code generated), confidence, or relevance of GPT’s solutions, enabling users to make more informed decisions and ensuring that developers maintain direction and avoid veering into unproductive tasks.
Moreover, researchers could explore the development of features that encourage reflective practices, such as visualizations that compare GPT’s suggestions with best practices or community standards. These features could help users critically evaluate AI-generated solutions rather than passively adopting them. Such studies should examine their impact on developer productivity, decision-making, and team dynamics over time. Longitudinal studies could provide valuable insights into how extended use of GPT affects user skill retention, adaptability, and overall collaboration in software engineering.
6 Threats to Validity
In this section, we discuss the threats to the internal and external validity of our study.
6.1 Construct Validity
Identifying GPT Use via Explicit Mentions: the use of a single keyword (”GPT”) to identify PRs that potentially involve the assistance from ChatGPT. This keyword-based method may result in false negatives, where GPT assistance was used but not explicitly mentioned, as well as false positives, where the term ”GPT” appears in a different context (e.g., in a discussion). Consequently, this introduces potential bias in the classification of PRs as GPT-assisted and may affect the accuracy of our findings regarding the actual impact of GPT use on the review process. While a semi-automated verification step helps mitigate this issue, future work could benefit from more robust detection methods, such as NLP-based intent analysis.
Definition of GPT as Assistance: A potential threat lies in how the GPT assistance is operationally defined. This study relies on heuristics to label PRs as GPT-assisted or non-assisted, but these heuristics may inadvertently misclassify instances, particularly when contributors do not explicitly document their use of GPT in PR titles, comments, or files. To mitigate this, manual inspections were performed to refine the heuristics and achieve high labeling precision (95%). However, some misclassification may persist.
Measurement of PR Phases: The study measures the duration of PR phases (e.g., review, waiting) using timestamps and events extracted from GitHub. While these metrics provide an objective basis for analysis, they may not fully capture qualitative aspects, such as the complexity of interactions or the nature of GPT’s contributions. In addition, inconsistencies in how timestamps and events are logged across projects could introduce variability. Efforts to mitigate these issues include cross-checking phase definitions and excluding PRs with missing or illogical events and timestamps.
Impact Attribution: Attributing observed reductions in PR phase times to GPT assistance alone is challenging. Other factors, such as task simplicity or team dynamics, could contribute to the observed effects. The use of a linear model helps account for some confounding variables (e.g., project age, number of commits), but residual confounding remains a potential threat. By addressing these threats, the study aims to provide a robust assessment of GPT’s influence on PR phases. However, future research could refine the constructs further, for example, by incorporating direct user feedback on their interactions with GPT or using more sophisticated techniques to distinguish GPT contributions from other factors.
Interactions with GPT: The nature of developer interactions with GPT may not be fully captured by simple time measurements. For example, the impact of GPT on developer behavior (e.g., whether it encourages more efficient coding practices or simply accelerates repetitive tasks) is not directly measured in this study. Future research could incorporate qualitative data, such as developer surveys or interviews, to better understand the cognitive and collaborative effects of GPT assistance on the code review process.
6.2 Internal Validity
Confounding Variables: The impact of GPT on PR resolution times may be influenced by factors other than GPT assistance, such as the complexity of the code changes, the experience of the developer, or the team’s workflow dynamics. While the study controls for some confounding variables through regression models, residual effects may still persist. Further steps, such as matching or propensity score analysis, could help control for additional confounders that were not included in the model.
Task Categorization: Categorizing the tasks for which GPT is used could be subjective and may vary depending on the developer’s interpretation of GPT’s assistance. If task categorization is inconsistent or unclear, it could skew the analysis and affect the conclusions drawn about the types of assistance most commonly used.
The Role of the Linear Model in Time Reduction: While the study finds a reduction in time for various phases of the code review process when GPT is used, it is important to ensure that this reduction in time is genuinely due to GPT’s assistance and not influenced by other factors (e.g., more efficient developers or less complex PRs). If time reduction is attributed to factors other than GPT, it could undermine the internal validity of the study.
6.3 External Validity
Platform-Specific Bias: This study analyzes GitHub data, which may have unique characteristics that limit the generalizability of findings to other platforms such as GitLab or Bitbucket. For instance, GitHub as a large open-source community and due to its widespread usage may introduce dynamics that are not presented in smaller or private repositories. Future research could include multiple platforms to better assess the broader applicability of these findings.
Repository Selection: The data included in this study were selected based on a threshold of 10 stars, indicating a relatively active user base. While this ensures that the results reflect commonly used, publicly available code, it also introduces selection bias. The results may not generalize well to less popular or private repositories, where developer practices may differ. Expanding the dataset to include a wider range of repositories, including those with lower visibility or fewer contributors, would help assess the external validity of the findings.
Task Variability: PRs vary widely in terms of content, including code size, task complexity, and the nature of the proposed changes. Although we attempted to control for these factors, the diversity in PRs introduces a level of variability that could influence the generalizability of the results. To address this, future studies could consider stratified sampling or clustering based on PR characteristics to assess the impact of GPT more accurately across different types of tasks.
7 Conclusion
This study highlights the potential of generative AI tools, particularly large language models (LLMs) such as ChatGPT, to enhance collaborative software development. Through an analysis of their application in GitHub pull requests (PRs), we observe substantial improvements across key phases of the PR workflow, most notably within the review phase, which exhibits a 33% reduction in review duration and an 87% decrease in waiting time for acceptance. These results underscore the capacity of GPT to improve code review efficiency and support diverse tasks, including code enhancement, bug resolution, and documentation generation, with the most significant gains occurring during the review process. While the contributions of GPT are clearly beneficial in certain phases, its limited influence during the submission and post-acceptance phases points to areas for further development and optimization. Future research should investigate how GPT can be more effectively integrated across the entire PR lifecycle and evaluate its impact on long-term productivity. As LLMs continue to advance, they are poised to become essential components of software engineering, promoting greater efficiency, collaboration, and code quality. By harnessing the capabilities of these tools, both practitioners and researchers can shape the evolution of collaborative development workflows and drive innovation in the field of software engineering.
Appendix A
In the following, we provide a detailed list of the patterns found in project name, PR title, PR body, and modified file to identify PRs using GPT as assistance, along with their definitions and intended matching purpose.
P1: Pattern for Detecting GPT-Related Identifiers. This regex identifies phrases containing ”chatgpt,” ”gpt,” or ”openai,” ensuring that they are preceded by descriptive words. It prevents ”chat” or ”gpt” from appearing immediately after the separator, enforcing a structured naming convention for AI-related terms. Example: project chatgpt_123.
[…]
P2: Pattern for Detecting GPT-Related Keywords in URL. Matches URLs containing ”chatgpt”,”gpt”, or ”openai” while ensuring a valid structure. It supports both HTTP and HTTPS protocols, optional ”www” subdomains, and common top-level domains (TLDs). The pattern also captures query parameters and paths where AI-related keywords appear, making it ideal for filtering relevant web links. Example:https://my-site.ai/chatgpt-version
[…]
P3:Pattern in Detecting with GPT-related Keywords in Path and URL. matches file pathsthat include specific AI-related keywords like ”chatgpt,” ”gpt,” or ”openai,” while also supporting common file path structures. It handles both forward and backslashes, valid path components, and file extensions. Example: /home/user/gpt_data.csv
[…]
|(chatgpt|gpt|openai))+\.[a-z]{2,3})
P4: Pattern in Detecting GPT Keywords in File Paths. matches strings containing the keywords ”chatgpt,” ”gpt,” or ”openai” within file paths or URLs. It handles optional spaces, separators (like hyphens and underscores), and supports valid URL components (letters, digits, and special characters). It’s designed to identify AI-related content in file paths, query parameters, or directories while allowing flexible formatting. Example: chatgpt@example.com, chatbot@openai.com
[…]
P5: Pattern for Matching ChatGPT References in Phrases . matches phrases like ”like in the chatgpt” or ”like the one from chatgpt” at the end of a string. It efficiently captures variations in the phrasing, ensuring that the string ends with a reference to ”chatgpt” while supporting both expressions.
[…]
P6: Pattern for Matching ”Spam” Followed by AI Keywords. matches the word ”spam” followed by a space and one of the AI-related keywords (”chatgpt,” ”gpt,” or ”openai”) at the end of a string. It is optimized for simplicity and flexibility, ensuring accurate detection of such phrases.
[…]
P7: Pattern for Matching AI Keywords and Code-Related Terms. matches strings containing AI-related keywords (like ”chatgpt,” ”gpt,” ”openai,” or ”azureai”) followed by code-related terms such as ”label,” ”function,” ”class,” ”model,” or ”service.” It handles both singular and plural forms efficiently, making it ideal for detecting AI and code-related concepts in text or code.
[…]
(label|class|functions?|constructors?|
calls?|formatter|demos?|sources?|
connectors?|plugins?|endpoints?|
models?|apis?|gateways?|services?))
P8: Pattern for Matching AI Keywords with Numeric Suffixes. matches occurrences of AI-related keywords (”gpt”, ”chatgpt”, ”openai”) followed by optional spaces, hyphens, or underscores and a numeric suffix (up to 6 digits or decimals). It is designed to efficiently identify AI keywords with numeric versions or identifiers, and works for both single and repeated instances in the string.
[…]
P9: Pattern for Matching GRUB GPT References. matches instances of ”grub_gpt” or ”GRUB_GPT” in strings, with support for optional spaces, hyphens, underscores, digits, and periods preceding the keywords. It identifies these specific references in various contexts, including file paths, configuration strings, or any similar patterns.
[…]
P10: Pattern for Matching AI Keywords at the End of a String. matches strings that end with one of the AI-related keywords (”chatgpt”, ”gpt”, or ”openai”), allowing for optional spaces, URL-safe characters, and hyphens before the keyword. It is ideal for identifying these specific keywords in paths, URLs, or other contexts where they appear at the end of a string.
[…]
(chatgpt|gpt|openai)\$
P11: Pattern for Matching Disk Types and Partition Formats. matches common disk types and partition formats, including ”disk guid,” ”mbr,” ”qemu,” ”kvm,” and ”gpt partition(s).” It is optimized for detecting variations in case and optional spaces or plurals.
(disk guid|mbr|MBR|qemu|QEMU|kvm|
[…]
P12: Pattern for Matching ChatGPT Requests with Action Verbs. matches requests directed to ”ChatGPT” or ”GPT” using various action verbs like ”say,” ”write,” ”suggest,” ”think,” etc., allowing for optional spaces or hyphens between ”chat” and ”gpt.” It’s designed to identify common user queries to AI models, including instructions and requests for explanations or suggestions.
[…]
(say|get|write|suggest|
tell|think|explain|
[…]
P13: Pattern for Matching OpenAI Chat Links. matches URLs from https://chat.openai.com, including optional /share/ paths and any query parameters. It efficiently captures both the base URL and potential query strings, ensuring flexible matching of OpenAI chat links.
https:\/\/chat\.openai\.com(\/share\/?)
[…]
P14: Pattern for Matching Examples Referencing GPT and Related Models. matches phrases like ”some example by GPT,” ”some example by ChatGPT,” or similar phrases involving AI models like OpenAI, Chat-GPT, or Coderabbit. It captures variations in spacing, hyphenation, and model names while ensuring flexibility in the format of such references.
[…]
(gpt|chatgpt|openai|chat\-gpt|
[…]
P15: Pattern for Matching GPT Versions Followed by ”Flawless”. matches strings like ”gpt-123 flawless,” where ”gpt-” is followed by numeric values and any combination of alphanumeric characters and spaces, concluding with the word ”flawless.”
[…]
P16: Pattern for Matching GPT Requests to ”Make Some Suggestion”. matches strings like ”gpt-123 make some suggestion,” where ”gpt-” is followed by digits and any alphanumeric characters or spaces, with the phrase ”make some suggestion” at the end. It accounts for optional spaces in the phrase for flexibility.
[…]
P17: Pattern for Matching GPT Requests to ”Walk Me”. matches the phrase ”gpt walk me,” allowing for optional spaces between ”gpt,” ”walk,” and ”me,” making it flexible for user input variations.
[…]
P18: Pattern for Matching Actions Performed with GPT and Related Models. captures actions like ”ask,” ”create,” ”generate,” and more, followed by prepositions like ”by,” ”with,” or ”from,” and then identifies associations with GPT and related models (e.g., ”gpt,” ”chatgpt,” ”openai,” etc.). It’s ideal for recognizing specific action-based queries related to AI.
[…]
generate|document|investigate|advise|
suggest|modify|solve|support|update|
summary|response|change|power|polish|
utilize|comment|write|improve|collaborate|
[…]
Author Contributions
Antonio Collante: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing Original Draft, Writing - Review and Editing, Visualization Samuel Abedu: Conceptualization, Software, Data Curation, Writing - Review and Editing SayedHassan Khatoonabadi: Conceptualization, Methodology, Validation, Writing - Review and Editing, Project Administration Ahmad Abdellatif: Conceptualization, Validation, Writing - Review and Editing Ebube Alor: Data Curation Emad Shihab: Conceptualization, Supervision, Project Administration, Funding Acquisition
Acknowledgements
This work was supported by the NSERC CREATE grant number 555406, 2021.